OpenStudy (anonymous):

determine tangent line of the curv at (1,1) 4x^2+4xy+4y^2

OpenStudy (anonymous):

Isolate for y. Find the derivative. Substitue x (1) into your derivative to get your slope. y = mx + b m = slope Sub your point (x,y) in to find b.

OpenStudy (anonymous):

would the derivative b= 8x+4x+y'+8y

OpenStudy (anonymous):

8x + 4x + 4y + 8y = 12x + 12y therefore y = -x

OpenStudy (anonymous):

(4x)(y) is the product rule for finding the derivative. The first x the deriv of the second + the second x the deriv of the first.

OpenStudy (anonymous):

y = -1 (slope) therefore y = -1x + b 1 = -1(1) + b b = 2 therefore y = -x + 2 is your equation of the tangent.

OpenStudy (anonymous):

ok can you explain the (4x)(y) part

OpenStudy (anonymous):

because the book says the derivative is 4(2x+y)